### v1

# ShortcutProbe: Probing Prediction Shortcuts for Learning Robust Models

Abstract
Deep learning models often achieve high performance by inadvertently learning spurious correlations between targets and non-essential features. For example, an image classifier may identify an object via its background that spuriously correlates with it. This prediction behavior, known as spurious bias, severely degrades model performance on data that lacks the learned spurious correlations. Existing methods on spurious bias mitigation typically require a variety of data groups with spurious correlation annotations called group labels. However, group labels require costly human annotations and often fail to capture subtle spurious biases such as relying on specific pixels for predictions. In this paper, we propose a novel post hoc spurious bias mitigation framework without requiring group labels. Our framework, termed ShortcutProbe, identifies prediction shortcuts that reflect potential non-robustness in predictions in a given model’s latent space. The model is then retrained to be invariant to the identified prediction shortcuts for improved robustness. We theoretically analyze the effectiveness of the framework and empirically demonstrate that it is an efficient and practical tool for improving a model’s robustness to spurious bias on diverse datasets.
1 Introduction
Deep learning models have shown remarkable performance across domains, but this success is often achieved by exploiting spurious correlations  Sagawa et al. (2020); Liu et al. (2021); Yang et al. (2023); LaBonte et al. (2024); Ye et al. (2024); Zheng et al. (2024a) between spurious attributes or shortcut features Geirhos et al. (2020) and targets. For example, models have been found to use correlations between textures and image classes Geirhos et al. (2019) for object recognition instead of focusing on defining features of objects. This issue becomes even more problematic in high-stakes domains like healthcare. For instance, models predicting pneumonia were shown to rely on correlations between metal tokens in chest X-ray scans from different hospitals and the disease’s detection outcomes Zech et al. (2018), rather than the pathological features of pneumonia itself. The tendency of using spurious correlations is referred to as spurious bias. Models with spurious bias often fail to generalize on data groups lacking the learned spurious correlations, leading to significant performance degradation and non-robust behaviors across different data groups. This robustness issue can have severe social consequences, especially in critical applications.
Mitigating spurious bias is crucial for robust generalization across data groups with varying spurious correlations. Existing methods on mitigating spurious bias Sagawa et al. (2020); Kirichenko et al. (2023) rely on group labels. Group labels represent spurious correlations with class labels and spurious attributes. For example, (waterbirds, water) Sagawa et al. (2020) represents a spurious correlation between waterbirds and water backgrounds in the images of waterbirds with water backgrounds. Using group labels specifies explicitly the spurious correlations that a model should avoid. However, obtaining group labels requires expert knowledge and labor-intensive annotation efforts. Moreover, group labels fail to capture subtle spurious biases, such as using certain pixels in images for predictions.
In this paper, we propose a post hoc approach that can automatically detect and mitigate potential spurious biases in a model rather than relying on group labels. Our key innovation is reframing the task of identifying and mitigating spurious biases as detecting and leveraging prediction shortcuts in the model’s latent space. Prediction shortcuts are latent features derived from input embeddings and predominantly contribute to producing the same prediction outcome across different classes. In essence, prediction shortcuts represent non-defining features of certain classes that the model heavily uses for predictions. By operating in the model’s latent space, our approach leverages the expressiveness of latent embeddings, enabling direct identification of spurious biases across diverse input formats without requiring group labels.
We present our post hoc approach as a novel framework called ShortcutProbe, which first identifies prediction shortcuts in a given model and leverages them to guide model retraining for spurious bias mitigation. ShortcutProbe utilizes a probe set without group labels, typically containing a diverse mix of features, to uncover potential prediction shortcuts. These shortcuts are identified as latent features extracted from sample embeddings belonging to different classes but producing the same prediction outcome. By optimizing these features to maximize the model’s confidence in their corresponding predictions, ShortcutProbe effectively encodes spurious attributes in non-generalizable prediction shortcuts that the model overly relies on for predictions.
With the identified prediction shortcuts, ShortcutProbe mitigates spurious biases by retraining the model to be invariant to these shortcuts, as they are irrelevant to true prediction targets. This invariance is achieved by applying regularization during retraining, which ensures that the identified prediction shortcuts no longer contribute to the model’s predictions of the true targets.
We theoretically demonstrate that when the spurious attributes in the training data are new to the model as reflected by the high prediction loss, the tendency of using spurious attributes for predictions is high after training on the data. Our method aims to revert the process of learning spurious attributes by retraining the model so that the learned spurious attributes induce high prediction losses, effectively unlearning the spurious attributes and reducing the influence of spurious correlations.
In summary, our contributions are as follows:
  • We introduce ShortcutProbe, a novel post hoc framework for mitigating spurious bias without requiring group labels. ShortcutProbe identifies prediction shortcuts and leverages them as a form of regularization for training robust models.
  • We provide a theoretical analysis revealing that our spurious bias mitigation approach effectively unlearns spurious attributes.
  • Through extensive experiments, we show that our method successfully trains models robust to spurious biases without prior knowledge about these biases.
2 Related Works
Shortcuts and spurious bias.
Shortcuts are decision rules that perform well on standard benchmarks but fail to transfer to more challenging testing conditions Geirhos et al. (2020), such as data with distributional shifts. Spurious correlations describe superficial associations between spurious (non-essential) features and targets in data, and they can be used by machine learning models as shortcuts Geirhos et al. (2019). This shortcut-learning phenomenon in machine learning models results in spurious bias, the tendency to use spurious correlations in data for predictions. Shortcut features are spurious attributes that are used by models in predictions. Accurately identifying shortcut features involves domain knowledge and model interpretation techniques. In our work, we circumvent this by identifying prediction shortcuts that represent potential shortcut features in a model’s latent space, making our method a general and useful method for spurious bias detection and mitigation.
Spurious bias mitigation with group labels.
Group labels in training data indicate the presence of spurious correlations across different subsets of the data. Exploiting group labels during training helps mitigate the reliance on the specified spurious correlations. Existing approaches for spurious bias mitigation utilize group labels to balance data distributions during training Cui et al. (2019); He and Garcia (2009); Byrd and Lipton (2019), to formulate a distributionally robust optimization objective Sagawa et al. (2019), or to progressively expand group-balanced training data Deng et al. (2023). Although these methods achieve remarkable success in spurious bias mitigation, the reliance on training group labels becomes a barrier in practice, as obtaining such labels often requires domain knowledge and labor-intensive annotation efforts. Our method does not require group labels during training; instead, it automatically detects prediction shortcuts that encode potential spurious biases the model has developed in training.
Reducing reliance on group labels.
To alleviate the dependency on training group labels, existing approaches propose inferring training group labels through various means, such as identifying misclassified samples Liu et al. (2021), clustering hidden representations Zhang et al. (2022), employing invariant learning techniques Creager et al. (2021), or training group label estimators using a few group-labeled samples Nam et al. (2022). Nevertheless, validation group labels are still needed to specify which biases to address. Recent works Zheng et al. (2024b, c) use vision-language models to infer group labels. Last-layer retraining methods Kirichenko et al. (2023); LaBonte et al. (2024) leverage group-balanced validation data to fine-tune the last layer of a model. While our approach also involves retraining the last layer, it does not require group labels. Instead, we introduce a novel strategy that utilizes identified prediction shortcuts during retraining, leading to improved model robustness.
3 Preliminary
A spurious correlation is the correlation between a spurious attribute present in the training samples and a prediction target. For example, the class waterbird and the attribute water background might form a spurious correlation in the images of waterbird, where some of them have water backgrounds, e.g., pond or river, and some do not. Spurious attributes are not truly predictive of the targets. A group label, e.g., (waterbird, water background), consists of a prediction target and a spurious attribute.
Figure 1: Illustration of ShortcutProbe. (a) The framework uses a set of probe data $\mathcal{D}_{\text{prob}}$ to identify prediction shortcuts by learning a shortcut detector to extract similar features from samples of different classes $i$ and $j$ that are all predicted as the same class $j$. Feature extractor $e_{\theta_{1}}$ and classifier $h_{\theta_{2}}$ are frozen during this stage. (b) ShortcutProbe then retrains the classifier with the probe data (the loss of the probe data $\mathcal{L}_{\text{ori}}$) while using the identified prediction shortcuts as regularization (the loss of the prediction shortcuts $\mathcal{L}_{\text{spu}}$).
Consider a model $f_{\theta}:\mathcal{X}\rightarrow\mathbb{R}^{|\mathcal{Y}|}$ with parameter $\theta$ trained on a training dataset $\mathcal{D}_{\text{tr}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ with $N$ sample-label pairs, where $x_{i}\in\mathcal{X}$ denotes a sample in the input set $\mathcal{X}$, $y_{i}\in\mathcal{Y}$ denotes a label in the label set $\mathcal{Y}$, and ||$\cdot$|| ⋅ | denotes the size of a set. The model $f_{\theta}=e_{\theta_{1}}\circ h_{\theta_{2}}$ can be considered as a feature extractor $e_{\theta_{1}}:\mathcal{X}\rightarrow\mathbb{R}^{D}$ followed by a classifier $h_{\theta_{2}}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{|\mathcal{Y}|}$, where $\theta=\theta_{1}\cup\theta_{2}$, $h_{\theta_{2}}$ is the last layer of the model, and $D$ denotes the number of dimensions.
Due to the existence of spurious attributes in $\mathcal{D}_{\text{tr}}$, the model can exploit them for predictions, such as recognizing waterbirds by detecting the existence of water backgrounds Sagawa et al. (2019). This presents a challenge: It is hard to determine whether a high-performing model is truly robust or simply “right for the wrong reasons”, i.e., relying on spurious attributes. Although models with spurious biases typically exhibit degraded performance when the learned spurious attributes are absent from input data, e.g., a waterbird on a land background, it remains challenging to identify the specific spurious attributes encoded by the model without group labels, which hinders the development of effective spurious bias mitigation strategies.
4 Methodology
4.1 Method Overview
We introduce ShortcutProbe, a post hoc framework that automatically detects and mitigates prediction shortcuts without requiring group labels to specify which spurious biases to address. The framework comprises two key steps: (1) Prediction shortcut detection, where a probe set is used to identify prediction shortcuts, and (2) Spurious bias mitigation, where the identified shortcuts are used in retraining the model to mitigate spurious biases in the model.
We provide an overview of our framework in Fig. 1. The process begins with a set of probe data $\mathcal{D}_{\text{prob}}$, which typically contains samples with various spurious attributes that reflect a model’s non-robustness to spurious correlations. These samples are mapped to the latent embedding space of the model $f_{\theta}$ through its feature extractor $e_{\theta_{1}}$, allowing us to model any prediction shortcuts the model might use for predictions. This strategy bypasses the need to explicitly define spurious correlations through group labels, a task that is especially challenging for subtle features, such as specific pixels in images. More concretely, as shown in Fig. 1(a), we first train a shortcut detector that extracts potential prediction shortcuts from samples of different classes but having the same prediction. The identified prediction shortcuts encode spurious attributes shared across classes and capture the model’s non-robustness across different data groups. Next, in Fig. 1(b), the model is retrained with $\mathcal{D}_{\text{prob}}$ to mitigate spurious biases by unlearning the identified prediction shortcuts.
4.2 Prediction Shortcut Detection
Given a model $f_{\theta}$ and a probe set $\mathcal{D}_{\text{prob}}$, we aim to detect prediction shortcuts by learning a shortcut detector $g_{\psi}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ to identify prediction shortcuts from input embeddings. Intuitively, prediction shortcuts can be identified from samples of different classes but having the same prediction outcome, an indication that similar features exist in these samples but are irrelevant to classes. In other words, prediction shortcuts can be shared among samples from different classes, necessitating a shared representational space to encode diverse prediction shortcuts. We formalize this intuition in the following definition.
Definition 1 (Prediction shortcuts).
Given input sample-label pairs $(x,y)$ and $(x^{\prime},y^{\prime})$, where $y\neq y^{\prime}$, a trained model $f_{\theta}=e_{\theta_{1}}\circ h_{\theta_{2}}$, sample embeddings $\mathbf{v}=e_{\theta_{1}}(x)$ and $\mathbf{v^{\prime}}=e_{\theta_{1}}(x^{\prime})$ for $x$ and $x^{\prime}$, respectively, a vector space $\mathcal{V}\subset\mathbb{R}^{D}$ spanned by $K$ base column vectors in $\mathbf{A}\in\mathbb{R}^{D\times K}$, prediction shortcuts $\mathbf{s}_{x}\in\mathcal{V}$ and $\mathbf{s}_{x^{\prime}}\in\mathcal{V}$ for the two samples satisfy the following conditions:
  • $\text{Pred}\big{(}h_{\theta_{2}}(e_{\theta_{1}}(x)\big{)}=y$, $\text{Pred}\big{(}h_{\theta_{2}}(\mathbf{s}_{x})\big{)}=y$, and
  • $\text{Pred}\big{(}h_{\theta_{2}}(e_{\theta_{1}}(x^{\prime})\big{)}=y$, $\text{Pred}\big{(}h_{\theta_{2}}(\mathbf{s}_{x^{\prime}})\big{)}=y$,
where $\text{Pred}(f_{\theta}(x))=\arg\max_{\mathcal{Y}}f_{\theta}(x)$, $\mathbf{s}_{x}=\mathbf{P}_{\mathbf{A}}\mathbf{v}$, $\mathbf{s}_{x^{\prime}}=\mathbf{P}_{\mathbf{A}}\mathbf{v}^{\prime}$, and
$\mathbf{P}_{\mathbf{A}}=\mathbf{A}(\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}^{T}$ (1)
is the projection matrix such that the prediction shortcut $\mathbf{s}_{x}$ is the best estimate of $\mathbf{v}$ in the vector space $\mathcal{V}$ in the sense that the distance $\|\mathbf{s}_{x}-\mathbf{v}\|_{2}^{2}$ is minimized.
In Definition 1, we define a prediction shortcut as a projection of a sample embedding. It exists in the vector space $\mathcal{V}$ shared by samples of different classes with $K$ degrees of freedom. Here, $K$ governs the complexity of the vector space representing prediction shortcuts. A smaller $K$ results in a less expressive vector space that may fail to adequately capture prediction shortcuts. Conversely, a larger $K$ provides greater flexibility in representing prediction shortcuts but may encode irrelevant information. The optimal value of $K$ depends on the complexity of the probe data; values that are too small or too large can impede learning and lead to suboptimal performance. We treat $K$ as a tunable hyperparameter.
By representing prediction shortcuts as vectors, we can in theory capture any spurious bias, even the intricate ones. For instance, a vector $\mathbf{s}$ might correspond to features of water backgrounds that are used to predict waterbirds in any image with water backgrounds, revealing a spurious bias in the model. Alternatively, $\mathbf{s}$ could represent a specific feature corresponding to certain pixels in input images, capturing the prediction shortcut based on low-level pixel values—an aspect that is challenging to articulate through group labels.
Learning the shortcut detector.
Based on the definition of the prediction shortcut, we design the shortcut detector $g_{\psi}$ as a function that implements the projection operation defined in Eq. (1) with the learnable parameter $\psi=\mathbf{A}\in\mathbb{R}^{D\times K}$, that is for a sample embedding $\mathbf{v}\in\mathbb{R}^{D}$, $g_{\psi}(\mathbf{v})=\mathbf{P}_{\mathbf{A}}\mathbf{v}$. Learning $g_{\psi}$ essentially learns a shared vector space spanned by $\mathbf{A}$ that could cover prediction shortcuts in samples in the probe set.
To effectively learn $g_{\psi}$, for each class $y$, we first collect samples from the probe set $\mathcal{D}_{\text{prob}}$ to formulate two non-overlapping sets $\mathcal{D}_{\text{cor}}^{y}$ and $\mathcal{D}_{\text{pre}}^{y}$, i.e.,
$\mathcal{D}_{\text{cor}}^{y}=\{(x,y)|(x,y)\in\mathcal{D}_{\text{prob}},\text{Pred}(f_{\theta}(x))=y\},$ (2)
and
$\mathcal{D}_{\text{pre}}^{y}=\{(x^{\prime},y^{\prime})|(x^{\prime},y^{\prime})\in\mathcal{D}_{\text{prob}},\text{Pred}(f_{\theta}(x^{\prime}))=y\neq y^{\prime}\},$ (3)
where $\mathcal{D}_{\text{cor}}^{y}$ and $\mathcal{D}_{\text{pre}}^{y}$ contain samples that are correctly and incorrectly predicted as $y$. The two sets together demonstrate a possibility that certain features shared across classes are incorrectly associated with prediction targets, allowing us to extract these features as potential prediction shortcuts.
Next, we propose the following objective to identify prediction shortcuts:
$\mathcal{L}_{\text{det}}=\mathop{\mathbb{E}}\limits_{y\in\mathcal{Y}}\mathop{\mathbb{E}}\limits_{(x,y)\in\mathcal{D}^{y}_{\text{cor}}\cup\mathcal{D}^{y}_{\text{pre}}}\ell\big{(}h_{\theta_{2}}(g_{\psi}(\mathbf{v})),y\big{)},$ (4)
where $\ell:\mathbb{R}^{|\mathcal{Y}|}\times\mathcal{Y}\rightarrow\mathbb{R}$ is the loss function, $\psi=\mathbf{A}$, and $\mathbf{v}=e_{\theta_{1}}(x)$. To ensure that prediction shortcuts are relevant to the given samples, we regularize the semantic similarity between $g_{\psi}(\mathbf{v})$ and $\mathbf{v}$ as follows,
$\mathcal{L}_{\text{reg}}=\mathop{\mathbb{E}}\limits_{y\in\mathcal{Y}}\mathop{\mathbb{E}}\limits_{(x,y)\in\mathcal{D}^{y}_{\text{cor}}\cup\mathcal{D}^{y}_{\text{pre}}}\|g_{\psi}(\mathbf{v})-\mathbf{v}\|_{2}^{2}$ (5)
The overall learning objective for $\psi$ is
$\psi^{*}=\arg\min_{\psi}\mathcal{L}_{\text{det}}+\eta\mathcal{L}_{\text{reg}},$ (6)
where $\eta>0$ represents the regularization strength. The objective in Eq. (6) aims to encode the properties of prediction shortcuts in Definition 1 into the shortcut detector $g_{\psi}$ while maintaining the relevance of the prediction shortcuts to input samples. Training $g_{\psi}$ is lightweight as there are only $DK$ learnable parameters. With $g_{\psi}$, we can identify multiple prediction shortcuts from samples in $\mathcal{D}_{\text{prob}}$.
4.3 Spurious Bias Mitigation
Mitigating spurious biases in a model requires that the spurious attributes captured during training are no longer predictive of the targets. Although identifying spurious attributes can be challenging, the shortcut detector introduced in the previous section identifies prediction shortcuts as potential spurious attributes utilized by the model, providing valuable guidance for addressing spurious biases.
In the following, we formulate an optimization objective that incorporates the above constraint to learn a robust model using the probe set $\mathcal{D}_{\text{prob}}$. First, a general requirement is that the trained model should produce correct and consistent predictions on training samples. To this end, for each class $y$, we sample from $\mathcal{D}^{y}_{\text{cor}}$ and $\mathcal{D}_{\text{mis}}^{y}$, where $\mathcal{D}^{y}_{\text{cor}}$ is the set of correctly predicted sample-label pairs defined in Eq. (2), and $\mathcal{D}_{\text{mis}}^{y}$ is defined as follows,
$\mathcal{D}_{\text{mis}}^{y}=\{(x^{\prime},y)|(x^{\prime},y)\in\mathcal{D}_{\text{prob}},\text{Pred}(f_{\theta}(x^{\prime}))\neq y\},$ (7)
representing misclassified sample-label pairs from the class $y$. Then, the training objective $\mathcal{L}_{\text{ori}}$ is as follows,
$\mathcal{L}_{\text{ori}}(\mathcal{D}_{\text{prob}};\theta)=\mathop{\mathbb{E}}\limits_{y\in\mathcal{Y}}\mathop{\mathbb{E}}\limits_{(x,y)\in\mathcal{D}^{y}_{\text{cor}}\cup\mathcal{D}^{y}_{\text{mis}}}\ell(f_{\theta}(x),y),$ (8)
which mitigates potential spurious biases by ensuring consistent predictions in $\mathcal{D}^{y}_{\text{mis}}$ and $\mathcal{D}^{y}_{\text{cor}}$.
Moreover, to ensure that the training targets at mitigating the spurious biases in the model, we further formulate a regularization term using the prediction shortcuts identified by our shortcut detector to guide the training process. Specifically, as the identified prediction shortcuts are not predictive of the targets, we aim to maximize the loss on the prediction shortcuts defined as follows,
$\mathcal{L}_{\text{spu}}(\mathcal{D}_{\text{prob}};\theta)=\mathop{\mathbb{E}}\limits_{y\in\mathcal{Y}}\mathop{\mathbb{E}}\limits_{(x,y)\in\mathcal{D}^{y}_{\text{cor}}\cup\mathcal{D}^{y}_{\text{mis}}}\ell(h_{\theta_{2}}(g_{\psi}(\mathbf{v})),y),$ (9)
where $\mathbf{v}=e_{\theta_{1}}(x)$.
We incorporate the above two loss terms into the overall training objective as follows,
$\theta_{2}^{*}=\arg\min_{\theta_{2}}\lambda\mathcal{L}_{\text{ori}}/\mathcal{L}_{\text{spu}},$ (10)
where $\lambda>0$ is the regularization strength. Here, we retrain only the final classification layer of the model while keeping the feature extractor frozen. This approach significantly reduces computational complexity and allows us to reuse the previously learned sample embeddings. Details of the training algorithm are provided in Appendix.
4.4 Choice of the Probe Set
The probe set plays a crucial role in both detecting prediction shortcuts and mitigating spurious biases. To achieve the full potential of ShortcutProbe, we use a held-out dataset—unseen by the model—to construct the probe set $\mathcal{D}_{\text{prob}}$. This choice is essential because the model may have memorized the training samples, making it difficult to identify prediction shortcuts based on discrepancies in predictions for samples of the same class. Moreover, we select samples with high predication confidence (measured by output entropy) from the held-out dataset to construct $\mathcal{D}_{\text{prob}}$ so that prediction shortcuts can be easily detected from these samples. Details of constructing $\mathcal{D}_{\text{prob}}$ and results on different choices of $\mathcal{D}_{\text{prob}}$ are provided in Section 5.
4.5 Theoretical Analysis
We theoretically demonstrate that minimizing the proposed objective $\mathcal{L}_{\text{ori}}/\mathcal{L}_{\text{reg}}$ effectively unlearns the spurious correlations between spurious attributes and their associated targets captured in the model. Without loss of generality, we analyze this in the context of a general linear regression setting. Consider an input sample $x\in\mathcal{X}$, a prediction target $y\in\mathcal{Y}$, and a spurious-only sample $\tilde{x}$ that lacks any defining features in $x$ related to $y$. Let $x_{1},\ldots,x_{N}$ denote $N$ training samples, $\varphi:\mathcal{X}\rightarrow\mathbb{R}^{D}$ be a generic feature map, and $J_{\mathbf{w}}(x)=\varphi(x)^{T}\mathbf{w}$ represent a generalized linear regression model with parameters $\mathbf{w}\in\mathbb{R}^{D}$.
We denote the correlation between the model output for a spurious-only sample $\tilde{x}$ and a prediction target $y$ as $\rho(J_{\mathbf{w}}(\tilde{x}),y)$. The following lemma Bombari and Mondelli (2024) gives an upper bound on the correlation.
Lemma 1.
The correlation between the model output for the spurious-only sample $\tilde{x}$ and the prediction target $y$ is upper bounded as follows:
$\rho(J_{\mathbf{w}}(\tilde{x}),y)\leq\gamma_{\varphi}\sigma_{\mathcal{Y}}\sqrt{\mathcal{R}_{\mathcal{X}}},$ (11)
where $\mathcal{R}_{\mathcal{X}}$ is the generalization error, $\sigma_{\mathcal{Y}}$ is the standard deviation of prediction targets, and
$\gamma_{\varphi}=\mathbb{E}_{\tilde{x},x}\Big{[}\frac{\varphi(\tilde{x})^{T}\mathbf{O}\varphi(x)}{\|\mathbf{O}\varphi(x)\|_{2}^{2}}\Big{]},$ (12)
where $\mathbf{O}=\mathbf{I}-\mathbf{V}^{T}(\mathbf{V}\mathbf{V}^{T})^{-1}\mathbf{V}$ is the orthogonal projection matrix, and $\mathbf{V}=[\varphi(x_{1}),\ldots,\varphi(x_{N})]^{T}\in\mathbb{R}^{N\times D}$ is the feature matrix.
Given that $\mathcal{R}_{\mathcal{X}}$ and $\sigma_{\mathcal{Y}}$ are independent of $\tilde{x}$, the spurious-only sample $\tilde{x}$ affects the correlation upper bound via the feature alignment term $\gamma_{\varphi}$ between the spurious attribute of $\tilde{x}$ and the original feature of $x$. We further interpret this term in the following proposition.
Proposition 1.
The feature alignment term $\gamma_{\varphi}$ is the ratio between the expected prediction error on the spurious sample $\tilde{x}$ and the expected prediction error on the original sample $x$:
$\gamma_{\varphi}$ $=\mathbb{E}_{\tilde{x},x}\Big{[}\frac{\varphi(\tilde{x})^{T}\mathbf{O}\varphi(x)}{\|\mathbf{O}\varphi(x)\|_{2}^{2}}\Big{]}.=\frac{\mathbb{E}_{\tilde{x}}[\|\mathbf{O}\varphi(\tilde{x})\|_{2}]}{\mathbb{E}_{x}[\|\mathbf{O}\varphi(x)\|_{2}]}$ (13)
The term $\|\mathbf{O}\varphi(x)\|_{2}$ in Eq. (13) denotes the error term for the sample $x$, while $\|\mathbf{O}\varphi(\tilde{x})\|_{2}$ denotes the error term for the spurious sample $\tilde{x}$. Since $x$ and $\tilde{x}$ are independent, the feature alignment term $\gamma_{\varphi}$ is the ratio between the loss on spurious-only samples and the loss on the original samples. We provide a detailed proof in Appendix.
The prediction shortcuts obtained by our shortcut detector approximate the spurious-only features. Thus, the loss $\mathcal{L}_{\text{spu}}$ approximates the nominator of the regularization term in Eq. (13). Moreover, $\mathcal{L}_{\text{ori}}$ approximates the denominator in Eq. (13). Note that Lemma 1 gives the upper bound of the spurious correlation before learning it. The objective in Eq. (4) trains the shortcut detector to learn the correlation by minimizing $\gamma_{\phi}$. In the mitigation step, the objective in Eq. (10) unlearns the correlation by maximizing $\gamma_{\phi}$.
5 Experiments
5.1 Datasets
Image datasets.
Waterbirds Sagawa et al. (2019) contains waterbird and landbird classes selected from the CUB dataset Welinder et al. (2010). The two bird classes are mixed with water and land backgrounds from the Places dataset Zhou et al. (2017). CelebA Liu et al. (2015) is a large-scale image dataset of celebrity faces. Images showing two hair colors, non-blond and blond, are spuriously correlated with gender. CheXpert Irvin et al. (2019) is a chest X-ray dataset containing six spurious attributes from the combination of race (White, Black, Other) and gender (Male, Female). Two diagnose results, i.e., “No Finding” (positive) and “Finding” (negative) are the labels. ImageNet-9 Ilyas et al. (2019) is a subset of ImageNet Deng et al. (2009) and contains nine super-classes. It is known to have correlations between object classes and image textures. We prepared the training and validation data as in Kim et al. (2022) and Bahng et al. (2020). ImageNet-A Hendrycks et al. (2021) is a dataset of real-world images, adversarially curated to test the limits of classifiers such as ResNet-50. While these images are from standard ImageNet classes Deng et al. (2009), they are often misclassified in multiple models. We used this dataset to test the robustness of a classifier after training it on ImageNet-9. NICO He et al. (2021) is designed for out-of-distribution image classification, simulating real-world scenarios where testing distributions differ from training ones. It labels images with both main concepts (e.g., cat) and contexts (e.g., at home). We used the Animal super-class in NICO and followed the setting in Bai et al. (2021); Tiwari and Shenoy (2023a) for data preparation.
Method Waterbirds CelebA Chexpert
WGA ($\uparrow$) Average ($\uparrow$) Gap ($\downarrow$) WGA ($\uparrow$) Average ($\uparrow$) Gap ($\downarrow$) WGA ($\uparrow$) Average ($\uparrow$) Gap ($\downarrow$)
ERM Vapnik (1999) 80.3±3.1 93.3±0.4 13.0 45.6±2.9 95.2±0.2 49.6 22.0±1.6 90.8±0.1 68.8
JTT Liu et al. (2021) 86.7±1.0 93.3±0.2 6.6 40.6±1.2 88.6±0.2 48.0 60.4±4.9 75.2±0.8 14.8
DFR Kirichenko et al. (2023) 90.3±2.1 95.0±1.3 4.7 72.2±2.0 92.9±0.1 20.7 72.7±1.5 78.7±0.4 6.0
AFR Qiu et al. (2023) 88.7±4.2 95.0±1.0 6.3 77.8±1.5 91.0±0.4 13.2 72.4±2.0 76.8±1.1 4.4
ShortcutProbe (Ours) 90.8±0.6 95.0±0.3 4.2 83.4±0.9 91.4±0.1 8.0 75.0±0.7 79.0±0.2 4.0
Table 1: Worst-group (WGA) and average accuracy (%) comparison with baseline methods on the Waterbirds, CelebA and Chexpert datasets. Best results are in boldface. All bias mitigation methods use the same half of the validation set.
Method MultiNLI CivilComments
WGA ($\uparrow$) Average ($\uparrow$) Gap ($\downarrow$) WGA ($\uparrow$) Average ($\uparrow$) Gap ($\downarrow$)
ERM Vapnik (1999) 67.0±0.4 82.2±0.2 15.2 58.5±1.3 92.2±0.1 33.7
JTT Liu et al. (2021) 71.6±0.8 80.7±0.4 9.1 68.3±0.9 89.0±0.3 20.7
DFR Kirichenko et al. (2023) 72.6±1.7 81.8±0.4 9.2 76.6±0.8 85.8±0.5 9.2
AFR Qiu et al. (2023) 66.6±0.3 82.2±0.2 15.6 74.6±5.1 84.7±2.5 10.1
ShortcutProbe (Ours) 74.3±0.7 82.6±0.3 8.3 79.9±0.6 88.5±0.2 8.6
Table 2: Worst-group (WGA) and average accuracy (%) comparison with baseline methods on the MultiNLI and Civilcomments datasets. Best results are in boldface. All bias mitigation methods use the same half of the validation set.
Text datasets.
MultiNLI Williams et al. (2017) is a text classification dataset with 3 classes: neutral, contradiction, and entailment, representing the natural language inference relationship between a premise and a hypothesis. The spurious feature is the presence of negation, which is highly correlated with the contradiction label. CivilComments Borkan et al. (2019) is a binary text classification dataset aimed at predicting whether a comment contains toxic language. Spurious features involve references to eight demographic identities: male, female, LGBTQ, Christian, Muslim, other religions, Black, and White.
5.2 Experimental Setup
Constructing the probe set.
From the chosen data source, such as the training or validation set, we sorted the samples within each class by their prediction losses and divided them into two equal halves: a high-loss set and a low-loss set. This approach approximates the incorrectly and correctly predicted samples, respectively, while ensuring that the incorrectly predicted set is non-empty, even when all samples are correctly classified. Within each set, we then selected the top $r$of samples with the highest prediction confidence (i.e., those with the lowest output entropy).
Training details.
We first trained a base model initialized with pretrained weights using empirical risk minimization (ERM) on the training dataset. Then, we retrained the model on half of the validation set using various bias mitigation methods. For our method, we first constructed the probe set using the same half of the validation set and used the probe set for shortcut detection and mitigation. The remaining half of the validation set is used for model selection and hyperparameter tuning. For experiments on the Waterbirds, CelebA, and Chexpert datasets, we used ResNet-50 as the backbone networks, and we used ResNet-18 on the ImageNet-9/A and NICO datasets to ensure a fair comparison with baseline methods. For text datasets, we use a pretrained BERT model Kenton and Toutanova (2019). We trained models with each method for three times using different random seeds and reported the averaged results as well as their standard deviations. Detailed training settings are provided in Appendix [^1] Code is available at https://github.com/gtzheng/ShortcutProbe..
Metrics.
Without group labels, we used the worst-class accuracy Yang et al. (2023) for model selection, which is defined as the worst per-class accuracy on an evaluation set. For performance evaluation on the Waterbirds, CelebA, Chexpert, MultiNLI, and Civilcomments datasets, we adopted the widely accepted metric, worst-group accuracy, which is the lowest accuracy among multiple groups of the test set with each group containing a certain spurious correlation. We also calculated the accuracy gap defined as the standard average accuracy minus the worst-group accuracy to measure the degree of a classifier’s spurious biases. A high worst-group accuracy with a low accuracy gap indicates that the classifier is robust to spurious biases and can fairly predict samples from different groups. We adopted average accuracy for the evaluations on the NICO, ImageNet-9, and ImageNet-A datasets as these datasets are specifically constructed to evaluate the robustness against distributional shifts.
5.3 Analysis of the Probe Set
Our method relies on the probe set for detecting prediction shortcuts and mitigating spurious bias. A good probe set can be used to effectively reveal and mitigate spurious biases in a model, such as those curated with group labels Sagawa et al. (2019). Here, we show that our method, ShortcutProbe, performs effectively without group labels when the probe set is carefully selected from readily available sources, such as the training data and held-out validation data.
To demonstrate, we constructed probe sets using the training data or half of the validation set (with the other half reserved for model selection) as described in Section 5.2. For each data source, we varied the proportion $r$ from 0.1 to 0.5. This adjustment created different versions of the probe set, ranging from those containing only samples with highly confident predictions (small $r$’s) to those including all samples from the selected data source ($r=0.5$).
Fig. 2(a) shows the performance of ShortcutProbe measured by worst-group accuracy (WGA) under different probe sets, while Fig. 2(b) presents the sizes of these probe sets. Compared to ERM models, we observe that reusing the training data for retraining results in minimal improvement on the Waterbirds dataset. This is because all the training data can be correctly predicted by the model, resulting in a probe set that is not informative to learn prediction shortcuts.
Our method demonstrates a significant improvement in robustness to spurious biases by leveraging a relatively small amount of held-out data compared to the size of the reused training data. Additionally, our approach benefits most from samples with high prediction confidence, i.e., when $r$ is small. However, as shown in Fig. 2(a), setting $r$ too small results in an insufficient number of training samples, leading to suboptimal WGA performance. In the following, unless otherwise specified, we use half of the validation set to construct the probe set and treat $r$ as a tunable hyperparameter.
Figure 2: Analyses on different probe sets constructed from the Waterbirds dataset. (a) Worst-group accuracy comparison between models trained with training data and half of the validation data. (b) Numbers of samples in respective probe sets.
5.4 Main Results
We focus on a challenging and practical setting where group labels are unavailable in both the training and validation data. This scenario requires detecting and mitigating spurious biases using only the data and models available in a standard ERM training setup. As baselines, we selected state-of-the-art last-layer retraining methods DFR Kirichenko et al. (2023) and AFR Qiu et al. (2023), as well as JTT Liu et al. (2021), which retrains the entire model. For DFR, which typically requires group labels, we used class labels instead. All baseline methods listed in Tables 1 and 2 utilized half of the validation data for training. Similarly, our method employed the same half of the validation data to construct a probe set, derived as a subset of this portion. The remaining half of the validation data was reserved for model selection.
As shown in Tables 1 and 2, our method achieves the highest WGA and the smallest accuracy gap between average accuracy and WGA, indicating its ability to strike a strong balance across different data groups. Methods that achieve high average accuracy, such as DFR on the CelebA dataset, prioritize learning spurious features in the probe set. Although maintaining good predictivity on average, DFR still suffer from the prediction shortcuts, as shown by its low WGA. Our method remains effective on larger backbone networks beyond ResNet-50, such as ResNet-152 and ViT (see Appendix). Unlike other baseline methods, our method uses only a portion of available data for training, highlighting the effectiveness of the probe set in detecting and mitigating spurious biases. Notably, when we applied the same probe set with baseline methods, this led to degraded performance in both WGA and accuracy gap, underscoring the unique advantages of our approach.
Method Accuracy ($\uparrow$)
ERM 75.9
REx Krueger et al. (2021) 74.3
Group DRO Sagawa et al. (2019) 77.6
JiGen Carlucci et al. (2019) 85.0
Mixup Zhang et al. (2017) 80.3
CNBB He et al. (2021) 78.2
DecAug Bai et al. (2021) 85.2
SIFER Tiwari and Shenoy (2023b) 86.2
ShortcutProbe (Ours) 90.5±0.6
Table 3: Comparison of average accuracy (%) on the NICO dataset.
We further tested the out-of-distribution generalization of our method on the NICO dataset. Images of each class in the test set are associated with an unseen context. Our method achieves the best classification accuracy without using group labels (Table 3), demonstrating its effectiveness in mitigating the reliance on contexts. We present additional results on the ImageNet-9 and ImageNet-A datasets in Appendix to demonstrate our method’s effectiveness in combating distributional shifts and the capability of achieving good tradeoff between in-distribution and out-of-distribution performance.
5.5 Ablation Studies
Figure 3: Analyses on how (a) prediction shortcuts as well as their regularization strength $\lambda$, (b) semantic regularization strength $\eta$, and (c) number of base vectors $K$ affect a model’s robustness to spurious biases. We report the worst-group accuracy on the CheXpert dataset.
Prediction shortcuts.
We evaluated the effectiveness of prediction shortcuts in Fig. 3(a). We began by using a randomly initialized shortcut detector to optimize the objective in Eq.(10). Additionally, we tested the performance without the spurious bias regularization term $\mathcal{L}_{\text{spu}}$, represented as $\lambda=0$ in Fig. 3(a), as well as for $\lambda$ values of 1, 10, and 50. Our results indicate that the model performs the worst when prediction shortcuts are not used as regularization. Interestingly, our method still demonstrates effectiveness even when the prediction shortcuts are random. As $\lambda$ increases, the regularization strength decreases. The model achieves optimal performance when an appropriately balanced $\lambda$ is selected.
Semantic regularization.
We analyzed the impact of the semantic regularization strength $\eta$ defined in Eq. (5). Fig. 3(b) shows that incorporating this regularization with a moderate value of $\eta$ enhances the model’s robustness.
Number of base vectors.
The number of base vectors, $K$, determines the representational capacity of the shortcut detector. A value of $K$ that is too small will limit the detector’s ability to identify spurious attributes effectively, while an excessively large $K$ may lead to overfitting on the probe data. Notably, as shown in Fig. 3(c), the optimal value of $K$ is 6, which coincides with the true number of spurious attributes in the CheXpert dataset.
6 Conclusion
In this work, we proposed a novel post hoc framework to mitigate spurious biases without requiring group labels. Our approach first learns a shortcut detector in the latent space of a given model via a diverse probe set. To mitigate spurious biases, we retrained the model to be invariant to detected prediction shortcuts using a novel regularized training objective. We theoretically demonstrated that this objective effectively unlearns the spurious correlations captured during training. Experiments confirmed that our method successfully mitigates spurious biases and enhances model robustness to distribution shifts. Future work may explore constructing a more diverse probe set to further enhance the detection and mitigation of spurious biases.
Acknowledgments
This work is supported in part by the US National Science Foundation under grants 2217071, 2213700, 2106913, 2008208.
Appendix A Appendix
A.1 Proof of Proposition 1
The proposition uses the result from Lemma 1 in our main paper, which is a restatement of Eq. (4.4) in Bombari and Mondelli [2024]. To prove the proposition, we first show that $\mathbf{O}\mathbf{O}=\mathbf{O}$. Note that $\mathbf{O}=\mathbf{I}-\mathbf{V}^{T}(\mathbf{V}\mathbf{V}^{T})^{-1}\mathbf{V}$, thus we have
$\mathbf{O}\mathbf{O}$ $=\big{(}\mathbf{I}-\mathbf{V}^{T}(\mathbf{V}\mathbf{V}^{T})^{-1}\mathbf{V}\big{)}\big{(}\mathbf{I}-\mathbf{V}^{T}(\mathbf{V}\mathbf{V}^{T})^{-1}\mathbf{V}\big{)}$
$=\mathbf{I}-2\mathbf{V}^{T}(\mathbf{V}\mathbf{V}^{T})^{-1}\mathbf{V}+\mathbf{V}^{T}(\mathbf{V}\mathbf{V}^{T})^{-1}\mathbf{V}$
$=\mathbf{I}-\mathbf{V}^{T}(\mathbf{V}\mathbf{V}^{T})^{-1}\mathbf{V}=\mathbf{O}.$ (14)
Next, we expand $\varphi(\tilde{x})^{T}\mathbf{O}\varphi(x)$ as follows
$\varphi(\tilde{x})^{T}\mathbf{O}\varphi(x)$ $=\varphi(\tilde{x})^{T}\mathbf{O}\mathbf{O}\varphi(x)$ (15)
$=(\mathbf{O}\varphi(\tilde{x}))^{T}(\mathbf{O}\varphi(x))$ (16)
$\leq\|\mathbf{O}\varphi(\tilde{x})\|_{2}\|\mathbf{O}\varphi(x)\|_{2},$ (17)
where Eq. (16) uses the fact that $\mathbf{O}^{T}=\mathbf{O}$, and (17) is the result of Cauchy–Schwarz inequality.
The inequality (17) holds in general. However, the equality actually holds in our setting. To show this, we need to prove that the vectors $\mathbf{O}\varphi(\tilde{x})$ and $\mathbf{O}\varphi(x)$ are independent. We first note that a spurious sample $\tilde{x}$ is independent of an original sample $x$. For example, in the Waterbirds dataset Sagawa et al. [2019], let $\tilde{x}$ represent an image showing only a water background, and $\tilde{x}$ is independent of $x$, as $\tilde{x}$ may be obtained by removing core objects from images of landbirds or waterbirds with waterbird backgrounds. Thus, the corresponding feature vectors $\varphi(x)$ and $\varphi(\tilde{x})$ are independent. If we assume that $\mathbf{O}\varphi(x)$ and $\mathbf{O}\varphi(\tilde{x})$ are dependent with $\mathbf{O}\varphi(x)=\eta\mathbf{O}\varphi(\tilde{x})$, where $\eta$ is a non-zero constant, then we have $\varphi(x)=\eta\mathbf{O}^{-1}\mathbf{O}\varphi(\tilde{x})=\eta\varphi(\tilde{x})$, which contradicts the fact that $\varphi(x)$ and $\varphi(\tilde{x})$ are independent. Therefore, $\mathbf{O}\varphi(x)$ and $\mathbf{O}\varphi(\tilde{x})$ are independent. Consequently, we have the following equality,
$\varphi(\tilde{x})^{T}\mathbf{O}\varphi(x)=\|\mathbf{O}\varphi(\tilde{x})\|_{2}\|\mathbf{O}\varphi(x)\|_{2}.$ (18)
Finally, we reinterpret the feature alignment $\gamma_{\varphi}$ as follows,
$\gamma_{\varphi}$ $=\mathbb{E}_{\tilde{x},x}\Big{[}\frac{\varphi(\tilde{x})^{T}\mathbf{O}\varphi(x)}{\|\mathbf{O}\varphi(x)\|_{2}^{2}}\Big{]}$ (19)
$=\mathbb{E}_{\tilde{x},x}\Big{[}\frac{\|\mathbf{O}\varphi(\tilde{x})\|_{2}\cdot\|\mathbf{O}\varphi(x)\|_{2}}{\|\mathbf{O}\varphi(x)\|_{2}\cdot\|\mathbf{O}\varphi(x)\|_{2}}\Big{]}$ (20)
$=\frac{\mathbb{E}_{\tilde{x}}[\|\mathbf{O}\varphi(\tilde{x})\|_{2}]}{\mathbb{E}_{x}[\|\mathbf{O}\varphi(x)\|_{2}]},$ (21)
where Eq. (21) results from the fact that the random variables $x$ and $\tilde{x}$ are independent.
A.2 Learning algorithm
We show the detailed training process of our proposed method, ShortcutProbe, in Algorithm 1. The algorithm is a two-step procedure. In the first step, we train a shortcut detector, and in the second step, we use the prediction shortcuts detected by the shortcut detector to mitigate spurious biases in the model.
Complexity analysis.
Given that the time complexity for obtaining $\mathcal{D}^{y}_{\text{cor}}$, $\mathcal{D}^{y}_{\text{pre}}$, and $\mathcal{D}^{y}_{\text{mis}}$ is $C_{\text{data}}$, the time complexity for each batch update during the shortcut detector training is $C_{\text{det}}$, and the time complexity for each batch update during classifier retraining is $C_{\text{ret}}$, the overall time complexity is $O(C_{\text{data}}+E_{1}N_{B}C_{\text{det}}+E_{2}N_{B}C_{\text{ret}})$.
Notably, the sets $\mathcal{D}^{y}_{\text{cor}}$, $\mathcal{D}^{y}_{\text{pre}}$, and $\mathcal{D}^{y}_{\text{mis}}$ can be precomputed before training and need to be constructed only once, allowing $C_{\text{data}}$ to be omitted once these sets are available. Additionally, $C_{\text{det}}$ and $C_{\text{ret}}$ are typically very small due to the lightweight design of the shortcut detector and the retraining process, which only involves the model’s final linear layer. Consequently, ShortcutProbe is highly computation-efficient. We provide a run-time comparison between different debiasing methods in Table 4 below.
JTT DFR AFR ShortcutProbe
1440 162 230 210
Table 4: Training time (s) comparison on the Waterbirds dataset.
Algorithm 1 ShortcutProbe
$\mathbf{Input}$: Probe set $\mathcal{D}_{\text{prob}}$, parameters of an ERM-trained model $\theta$ including $\theta_{1}$ of the feature extractor and $\theta_{2}$ of the classifier, parameters of the shortcut detector $\psi=\mathbf{A}$ with $K$ base vectors, batch size $B$, number of batches per epoch $N_{B}$, number of training epochs for the first step $E_{1}$, learning rate $\alpha$ used in the first step, number of training epochs for the second step $E_{2}$, learning rate $\beta$ used in the second step, regularization strengths $\eta$ and $\lambda$.
$\mathbf{Output}$: the classifier’s weights $\theta_{2}$
1:  Obtain $\mathcal{D}^{y}_{\text{cor}}$, $\mathcal{D}^{y}_{\text{pre}}$, and $\mathcal{D}^{y}_{\text{mis}}$ for each class $y$ from $\mathcal{D}_{\text{prob}}$ using Eq. (2) and Eq. (3), respectively
2:  //Learn shortcut detector
3:  for $e=1,\ldots,E_{1}$ do
4:     for $b=1,\ldots,N_{B}$ do
5:        Sample $\mathcal{B}^{y}_{\text{cor}}\subset\mathcal{D}^{y}_{\text{cor}}$ and $\mathcal{B}^{y}_{\text{pre}}\subset\mathcal{D}^{y}_{\text{pre}}$, $\forall y\in\mathcal{Y}$, with $|\mathcal{B}^{y}_{\text{cor}}|=|\mathcal{B}^{y}_{\text{pre}}|$ and $\sum_{y\in\mathcal{Y}}(|\mathcal{B}^{y}_{\text{pre}}|+|\mathcal{B}^{y}_{\text{cor}}|)=B$
6:        Calculate $\psi=\psi-\alpha\nabla_{\psi}(\mathcal{L}_{\text{det}}+\eta\mathcal{L}_{\text{reg}})$ using $\mathcal{B}^{y}_{\text{cor}}$ and $\mathcal{B}^{y}_{\text{pre}}$
7:     end for
8:  end for
9:  //Mitigate spurious biases
10:  for $e=1,\cdots,E_{2}$ do
11:     for $b=1,\cdots,N_{B}$ do
12:        Sample $\mathcal{B}^{y}_{\text{cor}}\subset\mathcal{D}^{y}_{\text{cor}}$ and $\mathcal{B}^{y}_{\text{mis}}\subset\mathcal{D}^{y}_{\text{mis}}$, $\forall y\in\mathcal{Y}$, with $|\mathcal{B}^{y}_{\text{cor}}|=|\mathcal{B}^{y}_{\text{mis}}|$ and $\sum_{y\in\mathcal{Y}}(|\mathcal{B}^{y}_{\text{mis}}|+|\mathcal{B}^{y}_{\text{cor}}|)=B$
13:        Calculate $\theta_{2}=\theta_{2}-\beta\nabla_{\theta_{2}}\lambda\mathcal{L}_{\text{ori}}/\mathcal{L}_{\text{spu}}$ using $\mathcal{B}^{y}_{\text{cor}}$ and $\mathcal{B}^{y}_{\text{mis}}$
14:     end for
15:  end for
16:  return  $\theta_{2}$
A.3 Datasets
[…]
Train Val Test
[…]
ImageNet-9 Ilyas et al. [2019] 9 N/A 54,600 2,100 N/A
ImageNet-A Hendrycks et al. [2021] 9 N/A N/A N/A 1087
[…]
Class Contexts
Validation Test
dog running in_street
cat on_tree in_street
bear on_tree white
bird on_shoulder in_hand
cow spotted standing
elephant in_circus in street
horse running in_street
monkey climbing sitting
rat running in_hole
sheep at_sunset on_road
Table 6: Classes and their associated contexts in the NICO datasets. Contexts not shown in the table are used in the training set.
NICO He et al. [2021] is a real-world dataset for evaluating a method’s out-of-distribution generalization performance. NICO provides context labels as spurious attributes. We used its Animal subset containing 10 object classes and 33 context labels. Following the setting in Bai et al. [2021]; Tiwari and Shenoy [2023b], we split the dataset into training, validation, and test sets with each set having unique contexts. Table 6 gives the allocation of the contexts for the 10 classes.
ImageNet-9 Bahng et al. [2020] is a subset of ImageNet, and ImageNet-A contains real-world images that are challenging to the image classifiers trained on standard ImageNet. Both datasets do not have clear group partitions specified by the class and attribute associations. ImageNet-9 has 9 super-classes, i.e., Dog, Cat, Frog, Turtle, Bird, Primate, Fish, Crab, Insect, obtained by merging similar classes from ImageNet. We extract images of the 9 super-classes from the ImageNet-A dataset and use these images for testing.
The CheXpert dataset Irvin et al. [2019] is a chest X-ray dataset from the Stanford University Medical center. There are six spurious attributes in the dataset, each of them is the combination of race (White, Black, Other) and gender (Male, Female). Two diagnose results, i.e., “No Finding” (positive) and “Finding” (negative) are the labels.
A.4 Training Details
ERM training.
This step trains ERM models which serve as the base models used in our framework for detecting prediction shortcuts and mitigating spurious biases. The training hyperparameters as well as the optimizer and learning rate scheduler used for each dataset are given in Table 7. For vision models, we initialized them with ImageNet-pretrained weights. For text models, we initialized them with weights pretrained on Book Corpus and English Wikipedia data.
Dataset Batch size Epochs Initial learning rate Weight decay Learning rate scheduler Optimizer
Waterbirds 32 100 0.003 0.0001 CosineAnnealing SGD
CelebA 128 20 0.003 0.0001 CosineAnnealing SGD
CheXpert 128 20 0.003 0.0001 CosineAnnealing SGD
MultiNLI 16 10 0.00001 0.0001 Linear AdamW
CivilComments 16 10 0.00001 0.0001 Linear AdamW
NICO 128 100 0.003 0.0001 CosineAnnealing SGD
ImageNet-9 128 100 0.003 0.0001 CosineAnnealing SGD
Table 7: Training settings for training ERM models on different datasets.
Training ShortcutProbe models.
We provide hyperparameter settings for the experiments on the Waterbirds, CelebA, CheXpert, MultiNLI, CivilComments, ImageNet-9, and NICO datasets in Table 8. We used an SGD optimizer with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$ in training the shortcut detector and retraining the classifier. We chose $K$ from $\{2,4,6,8\}$, $\eta$ from $\{0.1,1.0,5.0,10.0\}$, $\lambda$ from $\{0.1,1.0,5.0,10.0,50.0\}$, and $N_{B}$ from $\{50,100,200\}$, $\beta$ from $\{0.0001,0.0005,0.001,0.003,0.01\}$, and $r$ from $\{0.1,0.2,0.3,0.4,0.5\}$. We selected the best hyperparameters based on the performance on the whole validation set if the training data was used to construct the probe set or the remaining validation set if part of the validation set was used for the construction. The remaining hyperparameters were determined based on our empirical observations considering both dataset size and the convergence of training.
Training baseline models.
For JTT Liu et al. [2021], we combined the training data with half of the validation data to create a new training set for training JTT models. For DFR Kirichenko et al. [2023] and AFR Qiu et al. [2023], we applied these methods to the same ERM-trained model to ensure a fair comparison. We adhered to the hyperparameter settings recommended in the respective original papers.
Dataset $K$ $\eta$ $\lambda$ $E_{1}$ $E_{2}$ $B$ $N_{B}$ $\alpha$ $\beta$ $r$
Waterbirds 2 5.0 5.0 50 50 32 200 0.0001 0.001 0.3
CelebA 2 5.0 5.0 50 50 128 100 0.0001 0.001 0.1
CheXpert 6 10.0 50.0 50 50 128 50 0.0001 0.003 0.1
MultiNLI 2 5.0 5.0 50 50 128 100 0.0001 0.001 0.1
CivilComments 6 1.0 1.0 50 50 128 100 0.0001 0.003 0.1
NICO 8 1.0 1.0 50 50 128 200 0.0001 0.001 -
ImageNet-9 4 1.0 1.0 50 50 128 200 0.0001 0.001 -
Table 8: Hyperparameter settings for experiments on the seven datasets. $K$: number of base vectors; $\eta$: regularization strength for the semantic similarity constraint in Eq. (5); $\lambda$: regularization strength used in the training objective in Eq. (10); $E_{1}$: number of training epochs for learning the shortcut detector; $E_{2}$: number of training epochs for retraining the classifier; $B$: batch size; $N_{B}$: number of batches sampled in each epoch; $\alpha$: learning rate for learning the shortcut detector; $\beta$: learning rate for retraining the classifier; $r$: proportion of samples used to construct the probe set. When $r$ is not specified (“-”), it means using the training data to construct the probe set.
Appendix B Additional results
ImageNet-9 and ImageNet-A.
We presents performance comparison on the ImageNet-9 and ImageNet-A datasets in Table 10. The validation accuracy measures the in-distribution performance of a model, while the accuracy gap measures the performance drop from ImageNet-9 to ImageNet-A. Images in the ImageNet-A dataset represent distribution shifts from the training images in the ImageNet-9 dataset. Thus, the accuracy on the ImageNet-A dataset measures a model’s performance under distribution shifts. As shown in Table 10, our method achieves the best on the ImageNet-A dataset, demonstrating its robustness to distribution shifts. It also exhibits a good tradeoff between in-distribution and out-of-distribution performance by achieving the best accuracy gap.
ResNet-152 and ViT backbones.
Our method can be easily applied to larger backbone networks beyond ResNet-50, such as ResNet-152 and ViT. We evaluated our method with ResNet-152 and ViT-B/32 on three vision datasets and provide a performance comparison with baseline methods in Table 9 below. We observe that our method remains highly effective on large-scale models.
Backbone Method Waterbirds CelebA Chexpert
ResNet-152 ERM 17.4 60.6 18.6
DFR 30.3 68.3 66.0
AFR 31.4 70.7 63.8
Ours 34.7 80.0 68.3
ViT-B/32 ERM 69.5 52.2 20.9
DFR 87.6 63.0 73.4
AFR 86.6 79.5 64.5
Ours 88.0 83.7 75.1
Table 9: Comparison of worst-group accuracy (%) across last-layer retraining methods using ResNet-152 and ViT backbones.
Method ImageNet-9 ($\uparrow$) ImageNet-A ($\uparrow$) Acc. gap ($\downarrow$)
ERM 90.8±0.6 24.9±1.1 65.9
ReBias Bahng et al. [2020] 91.9±1.7 29.6±1.6 62.3
LfF Nam et al. [2020] 86.0 24.6 61.4
CaaM Wang et al. [2021] 95.7 32.8 62.9
SSL+ERM Kim et al. [2022] 94.2±0.1 34.2±0.5 60
LWBCKim et al. [2022] 94.0±0.2 36.0±0.5 58
SIFER Tiwari and Shenoy [2023a] 97.8±0.1 40.0±0.8 57.8
ShortcutProbe (Ours) 96.9±0.2 45.3±1.2 51.6
Table 10: Comparison of average accuracy (%) and accuracy gap (%) on the ImageNet-9 and ImageNet-A datasets.
Appendix C Qualitative Analysis of Learned Prediction Shortcuts
Figure 4: Visualization of the top-5 images that are most similar to the learned base vectors from the (a) Waterbirds and (b) CelebA datasets.
To qualitatively analyze the detected prediction shortcuts, we aim to interpret the learned base vectors in the matrix $\mathbf{A}$, as prediction shortcuts are obtained by linearly combining these vectors. To achieve this, for each base vector, we gave the top-5 images whose prediction shortcuts are most similar to the base vector. Specifically, for each learned base vector, we first calculated the embeddings of training samples, and following Eq. (1), we extracted prediction shortcuts in those samples as projected vectors by projecting the embeddings to the subspace spanned by the learned base vectors. We ranked the images based on the similarity of their prediction shortcuts to the base vector. A large similarity score signals a strong existence of the feature the base vector represents in the corresponding image.
As shown in Fig. 4(a), on the Waterbirds dataset, the two learned base vectors are most similar to images with land backgrounds and water backgrounds, respectively. In Fig. 4(b), the two learned base vectors are most similar to images of male celebrities and female celebrities, respectively. These results show that our shortcut detector can learn spurious attributes that well align with the biases in the datasets which models tend to capture during training.